tavon has a gift card for $175 that loses $4 for each​ 30-day period it is not used. He has another gift card for $155 that loses $3.50 for each​ 30-day period it is not used.


1
Expert's answer
2021-09-21T12:29:24-0400

"175-4\times x=155-3.5\times x\\175-155=4\times x- 3.5\times x\\20=0.5\times x\\x=40"

so the value of each gift card will be

"175-4\times x=155-3.5\times x"

now putting the value of x in above equation

"175-4\times 40=155-3.5\times 40\\175-155=155-140\\15=15"

so the value of each gift card will be $15.